3. Stock Valuation: You are the CEO of Under Armour, Inc. and are thinking about how to maximize your stock price. You believe your firm will pay dividends of $3 per share in exactly one year. Those dividends will grow at 25% for each of the next two years (e.g., D2 and D3 will grow at 25%) before settling down to a perpetual steady growth rate of 5% from year 4 onward. Under Armour’s current cost of equity is 18%.

a. What is your estimate of Under Armour’s current stock price?

b. After performing the analysis in part a), you reexamine your assumptions about Under Armour’s retention rate. You realize that your estimate of the firm’s perpetual growth rate (5%) is equal to the firm’s retention ratio (50%) and the expected return on new investment (10%). If you want to maximize the firm’s stock price, what is your optimal retention ratio in this scenario?

c. Suppose you identified a different investment opportunity with an expected return of 20%. Would your answer in part b) change? How and why?
Answer #1

The dividend D1 = $3

D2= $3*1.25= $3.75

D3= $3.75*1.25= $4.68

From D4 growth rateis constant at 5%.

Cost of equity =18%

Using Dividend discount model

a- Stock price= D1/(1+r) + D2/(1+r)^2 + D3/(1+r)^3+ D3/(R-g)

=44.14